# BAP v2: An Enhanced Task Framework for Instruction Following in Minecraft Dialogues

### v2

@@ 5.1.2 High-level Algorithm
    Architect Planning: First, A’s plan for the next steps is simulated, including which blocks to place or remove and their order if multiple blocks are involved. A reference block is chosen when necessary, enabling A to describe the new block(s) with appropriate spatial relations wrt the reference block. E.g., in Figure 11, A decides to place the red floating block, using the previously placed blue block as the reference.
@@
    Builder Position and Orientation: A "valid" position and orientation for B is sampled, ensuring a frame of reference for the utterances in the dialog segment. B is positioned at a suitable distance from the action area and oriented to face it while adhering to world constraints. E.g., in Figure 11, the structure $S^{t-1}$ is rotated according to B’s gaze (yaw angle only).  [^4] B’s position and pitch angle is omitted in our figures for clarity as noted in Section 2.2
@@
    Architect Instruction: An instruction from A is synthesized, incorporating necessary block details (e.g. color), spatial relations, and referential expressions, with template-based lexical and syntactic variations. E.g., in Figure 11, A instructs, "put a floating block …".
@@
    Optional Clarification Exchange: Some information may be omitted from the instruction with a small probability, allowing B to ask a clarification question. An appropriate exchange is then generated, with B asking and A providing the missing information, with template-based lexical and syntactic variations. E.g., in Figure 11, the block’s color is omitted, prompting B to ask, "what color?" and A to reply, "red".
@@
    Builder Actions and Low-Level Planning: B’s actions to add/remove blocks are then simulated, including any necessary low-level planning for deciding order of actions and accounting for temporary supporting blocks. This updates the built structure. E.g., in Figure 11, B places the red floating block to produce structure $S^{t}$, temporarily adding/removing support for the floating block.  [^5] Temporary supporting blocks are not shown in our figures as noted in Section 2.2.
  6. 6.
@@
    Optional Confirmation from Builder: Optionally, a confirmatory utterance from B, like "done," may be synthesized with a small probability.
Figure 11: A BAP item from one of our synthetic datasets
@@
We implement this general framework in three distinct ways, resulting in three different simulators. Each simulator uniquely instantiates steps 1, 3, and 4, along with independent target structure generation, while steps 2, 5, and 6 remain largely consistent across simulators. For brevity, we provide only a summary of each simulator in the following sections and provide more details in the appendix (Section 11).
5.1.3 Spatial Relations
Given the special emphasis on spatial relations in our dataset design, we briefly outline them here. We focus on the following types of spatial relations:
  1. 1.
@@ 5.1.3 Spatial Relations
    1. (a)
      1D: The target block lies in one of the rows, differing from the reference block in only one dimension. This results in 6 basic spatial relations: left, right, top, bottom, front, and behind (e.g., "to the right of …", "behind …").
@@
      2D: The target block lies in one of the quadrants, differing in exactly two dimensions. This results in 12 possible combinations of the 1D relations, such as top+right, bottom+left, etc. (e.g., "to the top right of …").
    3. (c)
      3D: The target block lies in one of the octants, differing in all three dimensions. This results in 8 possible combinations of the 1D relations, such as top+right+front (e.g., "one to the left, one above, and one block behind …").
  2. 2.
    Using B as a Spatial Anchor: Spatial relations can also be described using B as a spatial anchor directly (e.g., "going up and to the left of you"). These only apply when one needs to specify the direction in which to build a certain substructure/shape, e.g., a diagonal.
We will see concrete examples of these in the following sections.
5.2 Random target structures and dialogs
Dialogs
The dialogs correspond to random target structures generated dynamically during the dialog simulation (Section 5.1.2) which runs for a (bounded) random number of iterations. Blocks are placed/removed randomly, and one at a time in each iteration, with limited planning on A ’s part for block placement. As a result, the target structure is disordered, and A ’s instructions are fairly low-level, addressing only one block at a time. Most of them use spatial relations relative to arbitrary blocks identified by color and location (e.g., "to the right of the red" or "to the right of the bottom-most red block"). Figure 12 shows an example BAP item from this dataset. A identifies the orange block as a reference block and uses a 1D spatial relation "behind" to specify the target yellow block. We provide more details about the simulator in the appendix (Section 11).
Figure 12: A BAP item from the random structure-based dialogs
Target structures
Target structures are generated dynamically during the dialog simulation and correspond to the final built structure at the simulation’s end. Some global constraints are enforced:
  1. 1.
    The structure must have at least one block on the ground.
  2. 2.
Figure 13 provides an example.
Figure 13: A random target structure
5.3 Shape-based target structures and dialogs
Target structures in the MDC are not random; they are far more ordered (e.g., a dining table) and often composed of meaningful shapes/concepts (e.g., rows, towers, diagonals, planes, etc.). Architect instructions describe these structures at various levels of abstraction, from high-level references to the whole structure ("flower" or "bell tower"), mid-level references to sub-shapes ("row" or "plane"), down to low-level instructions referencing individual blocks. This observation motivates the need for similarly structured synthetic data as well. We first describe a simulator that can generate these structures (Section 5.3.1), and then outline two types of dialogs simulated based on these structures (Sections 5.3.2 and 5.3.3).
@@ 5.3.1 Shape-based structures
We define six unique elementary shapes: rows, diagonals, T-shapes, L-shapes, U-shapes, and planes. These shapes were chosen for their simplicity and occurrence in MDC dialogs and structures. (Formal definitions of the shapes are provided in the appendix (Section 11.4)). We design a simulator that generates target structures by combining multiple randomly sampled elementary shapes into a single composite structure. Sampling is performed with replacement, allowing multiple instances of the same shape type within a structure. For each shape instance, properties such as size, color, orientation, and exact 3D location are randomly sampled within specific constraints (e.g., minimum and maximum size bounds). Some global constraints are also enforced (similar to those for random structures in Section 5.2, with an added constraint for shapes):
  1. 1.
    The structure (i.e. at least one shape) must have at least one block on the ground.
  2. 2.
    The entire structure must be "connected", i.e., each block shares at least one face or edge with at least one other block, and each shape instance shares at least one block face or edge with at least one other shape instance.
The simulator is flexible, allowing parameters such as shape types, number of shapes per structure, etc. to be customized. Figure 14 illustrates two examples of such target structures.
Figure 14: Shaped-based target structures
5.3.2 Blocks-based dialogs for shape-based structures
Target structures
Each target structure consists of three shape instances, sampled from the full set of the six elementary shapes.
Dialogs
@@ 5.3.2 Blocks-based dialogs for shape-based structures
The dialog simulation (Section 5.1.2) takes the target structure as input and continues iteratively until the target structure is built. Construction proceeds one shape at a time. A ’s planning is involved and follows a heuristic that selects one or more blocks to place in each iteration until all shapes are completed. Consequently, A ’s instructions reference a single block or multiple ones collectively, with each shape decomposed into multiple instructions. Most of A ’s instructions include spatial relations relative to the last block placed by B (e.g., "now, place two red blocks behind the last block"). Figure 15 illustrates an example BAP item from this dataset. The first blue square has been built, and the second (yellow) shape is under construction. A is instructing B to continue building the latter by placing three more blocks, using a 1D spatial relation "on top of" relative to the last block. (In the previous step, A used a 2D spatial relation, "diagonally in front of and to the right of".) We provide more details about the simulator in the appendix (Section 11).
Figure 15: A BAP item from the blocks-based dialogs for shape-based structures
5.3.3 Shape-based dialogs for shape-based structures
Target structures
@@ 5.3.3 Shape-based dialogs for shape-based structures
For tractability in dialog simulation and to prevent overly complex utterances, each target structure is limited to two shape instances, sampled from only three elementary shapes: rows, diagonals, and planes.
Dialogs
@@
The dialog simulation (Section 5.1.2) takes the target structure as input and proceeds shape by shape, with one complete shape placed per iteration until the target is built. A ’s planning employs simple heuristics to determine the order of shapes and select a target block within each shape for defining spatial relations, which also serves as B ’s starting point for building the shape. A ’s instructions, therefore, describe entire shapes rather than individual blocks, resulting in higher-level, abstractive instructions. Since each target structure consists of two shapes, each placed in one go, a 1D/2D/3D spatial relation is only used to specify the position of the initial block of the second shape relative to the last block placed in the first shape. Additionally, A can also use spatial relations that use B as a spatial anchor (Section 5.1.3), for building either shape. B’s low-level planning also uses heuristics to determine the order of blocks within a shape, mimicking human Builders. Multiple orderings are possible, with some being more optimal or natural than others. E.g., 2D shapes like planes can be built either by column or by row. Additionally, B may also adopt a zigzag pattern, alternating the starting side for each row or column. Figure 16 illustrates a BAP item from this dataset. The first shape (a yellow column) has been built, and A is instructing B to build the second shape (a red row) in one go. A uses the spatial relation "going to the left of you" (using B as a spatial anchor) to indicate the direction for the row, along with the 3D spatial relation "one to the left, two underneath, and one block in front of" relative to the last block (the topmost yellow one), to indicate the starting position for the row. We provide more details about the simulator in the appendix (Section 11).
Figure 16: A BAP item from the shape-based dialogs for shape-based structures
6 Training with Synthetic Data

### v3

@@ 5.1 Our dialogue simulation framework
    Architect Planning (Step 1): The simulation begins by sampling the Architect’s plan for the next construction phase. This includes determining which blocks to place or remove and in what order. A reference block is chosen, when necessary, enabling A to describe the location of the new block(s) relative to the reference block in the subsequent dialogue $D^{t}$. In our running example, A decides to place a red floating block, using the previously placed blue block as a reference (diagonally from it at $(x_{1},y_{1},z_{1})$.
@@
    Adjusting the POV to a new Builder Position and Orientation (Step 2): As noted in Section 2.1.1, A and B can move around freely, but A views the structure from behind B when giving instructions. As a result, A instructions (including spatial relations, etc.) are grounded in B’s perspective. To emulate this, this step simulates B’s position and orientation $B^{t}$ (see Section 2.4 for a formal definition). A valid $B^{t}$ is sampled for B, establishing the frame of reference for the upcoming dialogue, $D^{t}$. $B^{t}$ is considered valid if B is at a suitable distance from the action area and oriented to face it, while respecting certain world constraints. E.g., Steps 2A and 2B show $B^{t}$ being chosen. In Step 2C, the grid is also rotated to align with B’s point-of-view and to view $S^{t-1}$ according to that POV – just for the reader’s visual clarity. Note that technically,
    Step 5.1.
    1 and
    Step 5.2.
    2 are independent and can be performed in any order.
@@
    Generating the Main Dialogue (Step 3):
    1. (a)
      Dialogue Style (Step 3A): The simulation first determines the dialogue style for the turn. It randomly chooses between a straightforward, complete instruction and, with much lower probability, an instruction that intentionally omits information to elicit a clarification question from the Builder.
    2. (b)
      Dialogue Generation (Step 3B): Next, the dialogue $D^{t}$ is synthesized using templates that incorporate lexical and syntactic variations. If a direct instruction was chosen, the utterance includes all necessary details (e.g., color, spatial relations (5.2.3), referential expressions). If a clarification was chosen, a key detail is omitted, and a corresponding question-answer exchange between B and A is generated. We show both possibilities for $D^{t}$ (omitting the block’s color for the clarification case).
@@
    Generating the Builder Actions (Step 4): The Builder’s actions ($A^{t}$) to execute the instruction are then simulated. This includes low-level planning to decide the precise order of actions and to account for temporary supporting blocks required for construction. These actions update the built structure, producing the new state, $S^{t}$. In our example, B places a red floating block to produce structure $S^{t}$, temporarily adding and removing a support block to do so. [^5] As noted in Section 2.2, temporary supporting blocks are not shown in our figures.
@@
    Generating an Optional Confirmation from Builder (Step 5): Finally, the turn may (with a low probability) conclude with a confirmatory utterance from B, such as “done!".
    The complete BAP item produced by this five-step process is shown in Figure 10 (bottom), which follows the standard visualization format used throughout this paper, so that structures like $S^{t-1}$ are rotated to align with the Builder’s POV $B^{t}$ (corresponding to the next turn and dialogue segment $D^{t}$ below it), and not with $B^{t-1}$, whereas Steps 2A-2C in the top of Figure 10 explicitly show the transition from the previous perspective ($B^{t-1}$) to the new one ($B^{t}$).
    We developed three distinct simulators based on this general framework. While
    Step 5.3.
    2,
    Step 5.4.
    4, and
    Step 5.5.
    5 are implemented largely consistently across all three, each simulator features a unique implementation of Architect Planning (
    Step 5.6.
    1) and Generating the Main Dialogue (
    Step 5.7.
@@
    3), as well as its own target structure generation. For brevity, we provide only a summary of each simulator in the following sections and provide more details for each in the appendix (Section 10).
    5.2 Generating Grounded Aspects of the Synthetic Dialogs
    We now briefly elucidate some key details of
    Step 5.8.
    2 and the spatial relations part of
    Step 5.9.
    3B, common across all three simulators.
    5.2.1 Sampling the Builder’s Position and Orientation
    Step 5.10.
    2
    For each turn in the simulation, we sample a new position and orientation for the Builder $\text{Pos}_{\textbf{B}}=(\text{loc}_{\textbf{B}},\phi_{\textbf{B}},\gamma_{\textbf{B}})$, where $\text{loc}_{\textbf{B}}=\langle x_{B},y_{B},z_{B}\rangle$ is the Builder’s location, $\phi_{\textbf{B}}$ is its pitch (vertical rotation), and $\gamma_{\textbf{B}}$ is its yaw (horizontal rotation) (Section 2.4). This sampling is grounded by the current state of the world, specifically the location of the reference block, $r_{p}$, which will be used as a spatial anchor in the Architect’s instruction.
    The process begins by identifying a set of all possible locations from which the Builder has an unobstructed view of the reference block. We then sample the Builder’s position, $\text{loc}_{\textbf{B}}$, uniformly from this feasible set. With the position established, we calculate the optimal yaw and pitch angles, $(\gamma_{\textbf{B}}^{*},\phi_{\textbf{B}}^{*})$, that would orient the Builder to look directly at the reference block from its new position:
    $\gamma_{\textbf{B}}^{*}$ $=\arctan\left(-\frac{r_{x}-x_{B}}{r_{z}-z_{B}}\right)$
    $\phi_{\textbf{B}}^{*}$ $=\arctan\left(\frac{y_{B}-r_{y}}{\sqrt{(x_{B}-r_{x})^{2}+(z_{B}-r_{z})^{2}}}\right)$
    To introduce natural variation and avoid perfectly robotic behavior, the Builder’s final orientation is sampled from a normal distribution centered on these optimal angles:
    $\gamma_{\textbf{B}}$ $\sim\mathcal{N}(\gamma_{\textbf{B}}^{*},h/2)$
    $\phi_{\textbf{B}}$ $\sim\mathcal{N}(\phi_{\textbf{B}}^{*},v/2)$
    where $h$ and $v$ are the Builder’s horizontal and vertical Field of View (FOV), respectively. We assume a “normal” game setting and 16:9 ratio of FOV, i.e., $h=102.4^{\circ}$ and $v=70^{\circ}$. [^6] https://minecraft.fandom.com/wiki/Options This ensures the Builder is generally facing the area of interest without being perfectly fixated on it.
    5.2.2 Grounding Spatial Relations in B’s Perspective (
    Step 5.12.
    3B)
    Computing perspective coordinates
    Once the Builder’s POV for the turn is established, we can generate spatial descriptions that are naturally grounded in this perspective. This is achieved by converting absolute world coordinates into ‘perspective’ coordinates (Jayannavar, Narayan-Chen, and Hockenmaier, 2020), which remap the 3D space relative to the Builder’s position and orientation. Any absolute location $\text{loc}=\langle x_{c},y_{c},z_{c}\rangle$ is transformed into perspective coordinates $\text{loc}^{\prime}=\langle x_{c}^{\prime},y_{c}^{\prime},z_{c}^{\prime}\rangle$ by first translating the origin to the Builder’s location $\text{loc}_{\textbf{B}}$, and then applying rotations to account for the Builder’s yaw ($\gamma_{\textbf{B}}$) and pitch ($\phi_{\textbf{B}}$): [^7] The rotation matrices for pitch ($P$) and yaw ($Y$) are defined as: $P\!=\!\left(\begin{smallmatrix}1&0&0\\ 0&\cos\phi_{\textbf{B}}&\sin\phi_{\textbf{B}}\\ 0&-\sin\phi_{\textbf{B}}&\cos\phi_{\textbf{B}}\end{smallmatrix}\right)$ and $Y\!=\!\left(\begin{smallmatrix}\cos\gamma_{\textbf{B}}&0&-\sin\gamma_{\textbf{B}}\\ 0&1&0\\ \sin\gamma_{\textbf{B}}&0&\cos\gamma_{\textbf{B}}\end{smallmatrix}\right)$.
    $\langle x_{c}^{\prime},y_{c}^{\prime},z_{c}^{\prime}\rangle=P\cdot Y\cdot\langle x_{c}-x_{B},y_{c}-y_{B},z_{c}-z_{B}\rangle$
    Computing spatial relations
    We leverage these perspective coordinates to generate phrases describing the position of the next block to be placed, $p$, relative to the reference block, $r$. We first convert both their absolute positions, $p_{p}$ and $r_{p}$, into perspective coordinates, yielding $p_{p}^{\prime}$ and $r_{p}^{\prime}$. The choice of directional words (e.g., "left/right" vs. "front/behind") is determined by comparing the difference between these new coordinates along the axes.
    For example, to determine the primary horizontal relation, we compare the magnitude of the difference along the Builder’s perspective x-axis (left/right) with the difference along the z-axis (front/behind). If $|p_{x}^{\prime}-r_{x}^{\prime}|>|p_{z}^{\prime}-r_{z}^{\prime}|$, the most salient direction is left or right. The specific term is chosen based on the sign of the difference: if $p_{x}^{\prime}>r_{x}^{\prime}$, the instruction would be "to the left of" the reference block. Conversely, if the difference along the z-axis is more significant, the instruction would use "in front of" or "behind". This method provides a deterministic way to generate simple and intuitive spatial language that is directly tied to what the Builder would be seeing from its simulated vantage point.
    5.2.3 Types of Spatial Relations
    We generate the following types of spatial relations:
    1. (a)
@@ 5.2.3 Types of Spatial Relations
      1. i.
        1D: The target block lies in one of the rows, differing from the reference block in only one dimension. This results in 6 basic spatial relations: left, right, top, bottom, front, and behind (e.g., “to the right of …", “behind …").
      2. ii.
        2D: The target block lies in one of the quadrants, differing in exactly two dimensions. This results in 12 possible combinations of the 1D relations, such as top+right, bottom+left, etc. (e.g., “to the top right of …").
      3. iii.
        3D: The target block lies in one of the octants, differing in all three dimensions. This results in 8 possible combinations of the 1D relations, such as top+right+front (e.g., “one to the left, one above, and one block behind …").
@@
      Using B as a Spatial Anchor: Spatial relations can also be described using B as a spatial anchor directly (e.g., “going up and to the left of you"). These only apply when one needs to specify the direction in which to build a certain substructure/shape, e.g., a diagonal.
    We will see concrete examples as we describe each dataset simulator in turn.
    5.3 Dataset $D_{\mathrm{r}}$: Dialogues for random target structures
    In this dataset, target structures are generated dynamically as a direct outcome of the dialogue simulation, corresponding to the final configuration of blocks at the end of the game. Dialogues are generated by running the simulation framework (Section 5) for a bounded, random number of turns. The key characteristics of this simulator emerge from its specific implementation of the Architect’s planning and dialogue generation steps. The Architect Planning (
    Step 5.14.
    1) process is intentionally minimalistic: at each turn, the Architect decides to place or remove a single block at a random valid location adjacent to the existing structure. This lack of sophisticated, high-level planning results in the dynamically generated, disordered target structures mentioned above, and directly influences the Generating the Main Dialogue (
    Step 5.15.
    3). Because the Architect considers only one block at a time, the resulting instructions are inherently low-level. They typically use 1D or 2D spatial relations that are anchored to arbitrary reference blocks, which are identified by their color or location (e.g., “behind" or “to the right of the red" or “to the right of the bottom-most red block", or, as in Figure 9 (top), “put a yellow block behind the orange one". Further details on this simulator are available in Appendix 10.
    5.4 Datasets $D_{\mathrm{sb}}$ and $D_{\mathrm{ss}}$: Dialogues for shape-based target structures
    Target structures in the MDC are not random; they are far more ordered (e.g., a dining table) and often composed of meaningful shapes/concepts (e.g., rows, towers, diagonals, planes, etc.). Architect instructions describe these structures at various levels of abstraction, from high-level references to the whole structure (“flower" or “bell tower"), mid-level references to sub-shapes (“row" or “plane"), down to low-level instructions referencing individual blocks. This observation motivates the need for similarly structured synthetic data as well. We first introduce a simulator that can generate such ordered and shape-based target structures (Section 5.4.1), and then outline two types of distinct dialogue simulators based on these structures (Sections 5.4.2 and 5.4.3).
@@ 5.4.1 Shape-based structures
    We define six unique elementary shapes: rows, diagonals, T-shapes, L-shapes, U-shapes, and planes. These shapes were chosen for their simplicity and occurrence in the MDC. (Formal definitions of the shapes are provided in the appendix (Section 10.5)). We then implemented a flexible target structure simulator that programmatically combines these elementary shapes into complex, composite structures. The simulator operates by randomly sampling multiple shape instances, allowing for repetition. For each instance, it further randomizes key properties such as size, color, orientation, and 3D location within specified constraints (e.g., minimum and maximum size bounds). The simulator is flexible, allowing parameters such as shape types, number of shapes per structure, etc. to be customized.
    5.4.2 Dataset $D_{\mathrm{sb}}$: Blocks-based dialogues for shape-based targets
    Each target structure consists of three shape instances sampled from the full set of six elementary shapes. Dialogues are generated by applying the simulation framework to the predefined shape-based structures until construction is complete. The simulation breaks down the construction of complex shapes into a sequence of simpler, block-level instructions. The Architect Planning (
    Step 5.16.
    1) is more sophisticated here. The Architect formulates a plan to build the target one shape instance at a time. Within each shape, it uses a heuristic to select one or more nearby blocks to place in each turn, ensuring a contiguous and logical build order. Consequently, the Generating the Main Dialogue (
    Step 5.17.
@@ 5.4.2 Dataset $D_{\mathrm{sb}}$: Blocks-based dialogues for shape-based targets
    3) produces low-level instructions that decompose each shape into multiple steps, referencing either a single block or a small group of blocks. A key linguistic feature of this dataset is that spatial relations are primarily anchored to the last block placed by B (e.g., “now, place two red blocks behind the last block"). Figure 9 (middle) provides examples. The first blue square has been built, and the second (yellow) shape is under construction. A instructs B to continue building the latter by adding blocks relative to the last one placed. A uses a 1D spatial relation “on top of" and the last block as the reference block. (In the previous step, A used a 2D spatial relation, “diagonally in front of and to the right of".) Further details on this simulator are available in Appendix 10.
    5.4.3 Dataset $D_{\mathrm{ss}}$: Shape-based dialogues for shape-based targets
@@
    This simulator generates dialogues at a higher level of abstraction than the block-based simulator described above, but, to ensure the dialogue simulation remains tractable and to avoid overly complex utterances, each target structure here is limited to two shape instances, sampled from only three elementary types: rows, diagonals, and planes. In Architect Planning (
    Step 5.18.
    1), the Architect plans to construct an entire shape instance within a single turn. The plan involves simple heuristics to select the order of the two shape instances and to identify a starting block for the second shape relative to the first. This high-level plan leads to more abstract, mid-level instructions during Generating the Main Dialogue (
    Step 5.19.
    3). Instead of referring to individual blocks, utterances describe whole shapes (e.g., “add a five block long red line…"). Spatial relations are used in two distinct ways: 1D, 2D, or 3D relations are used to specify the starting point of the second shape relative to the first, while relations that use B as a spatial anchor (Section 5.2.3) are used to define a shape’s orientation (e.g., “…going to the left of you"). This approach also simulates more complex Generating the Builder Actions (
    Step 5.20.
@@
    4). Given a high-level instruction to build a shape, the Builder must use its own heuristics to determine the actual sequence of block placements, mimicking how a human might build. E.g., 2D shapes like planes can be built either column-by-column or row-by-row. Additionally, B may also adopt a zigzag pattern, alternating the starting side for each row or column. Figure 9 (bottom) shows an example. The first shape (a yellow column) has been built, and A instructs B to build the second shape (an entire red line) in one go, using both types of spatial relations. A uses the spatial relation “going to the left of you" (using B as a spatial anchor) to indicate the direction for the row, along with the 3D spatial relation “one to the left, two underneath, and one block in front of" relative to the last block (the topmost yellow one), to indicate the starting position for the row. Further details on this simulator are available in Appendix 10.
    5.5 A Qualitative Comparison of the Datasets and MDC
[…]
    5.6 Related Work: Synthetic Data
    Synthetic data has proven beneficial for various tasks, including those mentioned in Section 2.8. VLN agents, constrained by limited human instruction data and diversity in training environments, often struggle with complex language grounding and spatial language understanding. Kamath et al. (2023) address this by using synthetic data, achieving SOTA performance on the RxR dataset (Ku et al., 2020), while Wang et al. (2023) are similar in spirit and achieve SOTA on the aforementioned CVDN and other benchmarks. Similarly, Kang et al. (2023) demonstrate strong performance gains in low-data regimes for Visual Dialogue Das et al. (2017) by leveraging synthetic data. Synthetic data has also been effective for dialogue systems, particularly those focused on text-only dialogues (Kim et al., 2022; Bao et al., 2023; Zhan et al., 2023). Zhan et al. (2024) is an example of extending this to multi-modal dialogue agents, demonstrating the effectiveness of synthetic visual descriptions in enhancing agents’ grounding capabilities. We have also seen synthetic data used to improve VLM spatial reasoning. Tang et al. (2024) do so by constructing simple images and fine-tuning over simple tasks such as 2D object localization, distance estimation, navigations, etc to improve spatial reasoning capabilities in 3D images.
    Synthesizing Task-Oriented Embodied Dialogues
    However, fewer efforts have addressed synthesizing complete task-oriented embodied dialogues (involving both utterances and environment actions). Padmakumar et al. (2023) were the first to design a framework to do so. They extend agenda-based dialogue simulation (Schatzmann and Young, 2009) to a multimodal embodied agent framework, and demonstrate the impact of the synthetic dialogues on the TEACh task (see Section 2.8 for a comparison of TEACh to MDC). Our work serves to provide another useful example in this relatively underexplored general domain, further advancing investigation of synthetic data generation for such complex tasks. Dan, Han, and Roth (2021) is another closely related work that demonstrates the potential of synthetic data in the Blocks World domain. They use simpler synthetic data for the benchmarks of Bisk, Yuret, and Marcu (2016), which require understanding single-shot instructions that transform one world state to another using simulated 3D blocks. Blocks are viewed from a fixed bird’s eye perspective, initialized randomly in the initial world state, and uniquely identifiable. The varying Builder perspective and lack of easily identifiable referents, along with the need to understand utterances in a dialogue context, make the BAP task a much more challenging problem. [^8] As noted in Section 2.1.1, unlike traditional Blocks World, Minecraft allows blocks to float (requiring non-monotonic action sequences where placement is followed by removal), or attach to any side of an existing block. Our work on synthetic data thus serves as a demonstration of the potential for richer, embodied task-oriented dialogue tasks that extend beyond simpler Blocks World settings.
    LLM-generated Synthetic Data
    While other domains have adopted LLM-generated synthetic data, embodied task-oriented dialogue has generally not followed suit. LLMs are able to generate diverse dialogues (Samarinas et al., 2024) or dialogue counterparts to an existing modality (Zhang et al., 2025a), but they struggle to generate the real-world environment and track state-changes within it, even on the simpler un-embodied task-oriented dialogue settings (Li et al., 2024b; Feng et al., 2023b; Kulkarni et al., 2024; Hudeček and Dusek, 2023). Generating dialogue from scratch (as Du et al. (2025) do) requires complex systems of planning and synthesis that are extremely domain-specific and must be evaluated by humans to verify coherence, flow, and accuracy. Our work, therefore, utilizes a programmatic simulation framework to overcome such limitations. This method guarantees that all generated data is accurately grounded in the game’s world state via highly nuanced geometric computations (Section 5.2), providing a reliable and systematic way to generate high-fidelity training data for complex, embodied tasks like BAP.
    6 Revisiting BAP Modeling: Training with Synthetic Data